Test the hypothesis that educational attainment does not depend on socio economic status for the following 100 persons in a particular community. ∝= 5%

Socioeconomic status/ Educ. Attainment is Poor , Finished College is 18 Did not finished college is 10. Middle Class: Finished College is 28, did not finished college is 25. Rich: finished college is 15 and did not finished college is 5


1
Expert's answer
2021-11-17T10:12:31-0500

Chi-Square test


"H_0:" sales and store location are independent

"H_a:" sales and store location are dependent


using online calculator www.socscistatistics.com, we get:


"\chi^2=\sum\frac{(O_i-E_i)^2}{E_i}"


where Oi is observed value,

Eis expected of cell:


"E_i=\frac{(row\ total)(column\ total)}{total\ number}"


"df=(rows-1)(columns-1)=(2-1)(3-1)=2"


The chi-square statistic is 3.2288. The p-value is 0.199013

Since p-value > 0.05, we accept the null hypothesis.

Educational attainment does not depend on socio economic status.